A parkour artist jumps from one rooftop to another. He jumps off the first rooftop with a velocity of 5m/s, angled at 30° to the horizontal. He stays airborne for 0.70 seconds before landing on the second rooftop.



What was the vertical velocity (𝑣⃗2𝑦) of the parkour artist when they reached the second rooftop?


1
Expert's answer
2021-11-04T10:17:59-0400

The time during which the body reached its maximum height


"t=v\sin30°/g=5\cdot\sin30°/9.81=0.255\ (s)"


"t_1=0.7-0.255=0.445\ (s)"


The vertical velocity


"v_y=gt=9.8\cdot0.445=4.37\ (m/s)"